Consider the square matrices

D (3x3)=
1 −1 1

3 2 2

3 -3 5
(i) Compute det(D). Write down det(D3), without computing D3
Answer #1

3 3-5 3 22 t I = 1 (10-(-67) +1 (15-6)+1 (-9-6 34(16)-1(90+1(15 士6キー5-15 -- Det = 10 10 10 34 (D):6101 キーーーー上 1000